# Dual membership certificatesJoonyeob Lee, Dávid Papp, Anita Varga

### v2

@@ Abstract
We discuss easily verifiable cone membership certificates, that is, certificates proving relations of the form $b\in\mathcal{K}$ for convex cones $\mathcal{K}$, that consist of vectors in the dual cone ${\mathcal{K}^{*}}$. In the standard theory of duality, vectors in $K^{*}$ are associated with separating hyperplanes and interpreted as certificates of non-membership in $K$. Complementing this, we present constructive certification schemes through which members of the dual cone can be interpreted as primal membership certificates. Every vector in the interior of $\mathcal{K}$ is assigned a full-dimensional cone of certificates, making the numerical computation of rigorous certificates easy, provided that the dual cone has an efficiently computable logarithmically homogeneous self-concordant barrier. Of particular interest are cones that are low-dimensional linear images of much higher dimensional cones. In the context of optimization (as opposed to feasibility) problems, these certificates can be used to easily compute, using a closed-form formula, exact primal feasible solutions from suitable dual feasible solutions, with the guarantee that the closer the dual solutions are to optimality, the closer to optimality are the computed primal solutions, too. We demonstrate that the new certification scheme is applicable to virtually every tractable subcone of nonnegative polynomials commonly used in polynomial optimization (such as SOS, SONC, SAGE and SDSOS polynomials, among others), facilitating the computation of rigorous nonnegativity certificates using numerical algorithms.
@@ 1 Introduction
for every $z\in\mathcal{K}\setminus\{0\}$ [13, Chapter 6]. This vector $y$ can be interpreted as a non-membership certificate, that is, a proof that $b$ does not belong to $\mathcal{K}$. By definition, $y$ is an element of the interior of the dual cone of $\mathcal{K}$,
@@
this gives rise to the standard duality theory of convex (conic) programming. In this paper we show that with the help of self-concordance theory, membership (rather than non-membership) in the primal cone can be similarly certified by simply exhibiting suitable vectors from the dual cone. These certificates are not unique. On the contrary, elements of the interior of $\mathcal{K}$ have a full-dimensional cone of membership certificates; this facilitates their computation using numerical methods and has important theoretical consequences.
Our first main contribution is that whenever ${\mathcal{K}^{*}}$ is a hyperbolicity cone, membership in $\mathcal{K}$ can be certified in a manner similar to the dual sums-of-squares certificates of [8] (we call these H-certificates in Definition 2.4), but the generalization does not go any further (Example 2.3). Then we present a new certification scheme (termed B-certificates) that is equally efficient for the same cones as H-certificates, but can also be applied to all cones $\mathcal{K}$ as long as ${\mathcal{K}^{*}}$ is the domain of an efficiently computable logarithmically homogeneous self-concordant barrier function (Theorem 2.6).
@@
In Section 3 we demonstrate that our dual membership certificates overcome both of these issues. First, because the cone of certificates for every vector is full-dimensional, numerically computed certificates are automatically exact, rigorous certificates; and since they are just elements of $\mathcal{K}_{A}^{*}$, they are not higher dimensional than the certified vector. Second, the certification schemes come with a reconstruction formula for cones of the form (1) that allows the efficient computation of a primal membership certificate $x\in\mathcal{K}$ that by definition satisfies $b=Ax$ exactly. (Theorem 3.2.) For most cones $\mathcal{K}$ that commonly arise in conic optimization problems, the formula automatically yields rational primal certificates $x$ (as long as $A$ is a rational matrix), which can also be verified efficiently in rational arithmetic.
In Section 6, we show how dual certificates can be used to compute, with a closed-form formula, exact primal feasible solutions of conic optimization problems in standard form from suitable dual feasible solutions, with a guarantee that the closer the dual solutions are to optimality, the closer the computed primal feasible solutions will be to optimality. With this formula, most short-step, path-following IPAs (including dual-only methods) can be “retrofitted” to compute primal-dual feasible pairs with a guaranteed low duality gap with minimal additional computational cost in each iteration.
In Section 7, we outline the application of these ideas in polynomial optimization. The most commonly used inner approximations of the cones of nonnegative polynomials (including sums-of-squares (SOS) polynomials, sums of nonnegative circuit (SONC) polynomials, sums of AM-GM exponentials (aka. SAGE functions), SDSOS polynomials and others) are all cones of the form (1) with an easily computable matrix $A\in\mathbb{Q}^{m\times n}$, usually with $m\ll n$, and a (possibly non-symmetric) cone $\mathcal{K}$ for which ${\mathcal{K}^{*}}$ admits a well-known logarithmically homogeneous self-concordant barrier function. Therefore, the problem of rigorously certifying the nonnegativity of a polynomial using either of these cones amounts to the exact solution of a feasibility problem in $\mathcal{K}_{A}$, as discussed above. When $\mathcal{K}$ is any of the aforementioned cones of polynomials, the dual certificates introduced in this paper can be interpreted as easily verifiable rigorous polynomial nonnegativity certificates that can be computed numerically without constructing an explicit “primal” representation (such as an explicit representation of an SOS polynomial as a sum of squared polynomials). Should such a representation be necessary, the reconstruction formula can be used to compute an exact representation in closed form using a numerically computed dual certificate.
@@
The results in this paper are broad generalizations of prior work of Davis and Papp [8] on weighted sums-of-squares (WSOS) cones, which are convex cones instrumental in global polynomial optimization. Membership in these cones can be cast as a (feasibility) semidefinite program (SDP) [30]; every feasible solution to this SDP is an obvious membership certificate. However, the number of variables in this SDP is at least an order of magnitude larger than the dimension of the cone (and in the presence of a large number of weights, the ratio of the dimensions of these two cones can be arbitrarily large). Instead, the certificates introduced in [8] (simply termed there as dual certificates) are elements of the dual WSOS cone, and are thus low-dimensional, and can be more efficiently computed using non-symmetric cone programming methods than solving the conventional SDP. The theory presented in the present paper is a generalization of these earlier results, and the algorithms discussed here are simultaneously more general and more efficient (owing to their lower iteration complexity) than the algorithms presented in the earlier work.
@@
    $f(x)\to\infty$ as $x\to\operatorname{bd}(\mathcal{C})$ and
@@
Proposition 1.2 below summarizes a few properties of self-concordant barrier functions used throughout the paper. The proofs of these statements can be found, for example, in [34, Chapters 2–3] and [24, Chapter 5].
@@
Suppose $f:\mathcal{C}^{\circ}\to\mathbb{R}$ is an SCB for some convex cone $\mathcal{C}\subseteq\mathbb{R}^{n}$, and let $g(x)$ and $H(x)$ denote its gradient and Hessian at $x\in\mathcal{C}^{\circ}$. Then the following hold:
@@
    The gradient map $g:\mathcal{C}^{\circ}\to\mathbb{R}^{n}$ is a bijection between $\mathcal{C}^{\circ}$ and $-(\mathcal{C}^{*})^{\circ}$. That is, for every $b\in(\mathcal{C}^{*})^{\circ}$ there exists a unique $y_{b}\in\mathcal{C}^{\circ}$ that satisfies $-g(y_{b})=b$.
@@
    If $A:\mathbb{R}^{m}\to\mathbb{R}^{n}$ is an injective linear operator and $\mathcal{A}\overset{\mathrm{def}}{=}\left\{x\in\mathbb{R}^{m}\,|\,A(x)\in\mathcal{C}^{\circ}\right\}$ is non-empty, then $x\mapsto f(A(x))$ is a $\nu$-LHSCB for $\mathcal{A}$.
@@ 2 Membership certificates from the dual cone
Throughout this section, we shall use the following assumptions and notational shortcuts: we assume that $\mathcal{K}$ is a proper convex cone whose dual cone ${\mathcal{K}^{*}}$ admits a $\nu$-LHSCB $f$, and we denote the gradient and Hessian of $f$ at $y\in{({\mathcal{K}^{*}})^{\circ}}$ by $g(y)$ and $H(y)$, respectively. Finally, we define
@@
The membership certification schemes introduced in this section are based on the following facts.
@@
Let $\mathcal{K}$ be a proper cone whose dual cone ${\mathcal{K}^{*}}$ is the domain of an SCB, and let $g$ and $H$ denote the gradient and Hessian of this barrier, respectively. Then for every $y\in{({\mathcal{K}^{*}})^{\circ}}$ and $z\in{\mathcal{K}^{*}}$,
@@
If ${\mathcal{K}^{*}}$ is the hyperbolicity cone of the hyperbolic polynomial $p$ and $f$ is the logarithmic barrier $-\ln(p(\cdot))$, then we also have
@@
The second claim is Güler’s result from [12, Thm. 6.2].
The assumption of hyperbolicity in the second half of Theorem 2.1 is necessary; the relation (5) does not hold for every cone.
@@
=$\mathcal{E}\overset{\mathrm{def}}{=}\operatorname{cl}\left(\left\{(x_{1},x_{2},x_{3})\in\mathbb{R}_{+}^{2}\times\mathbb{R}\,|\,x_{2}>0,x_{1}\geq x_{2}e^{x_{3}/x_{2}}\right\}\right),$
@@
see, e.g., [23, Eq. (4.5)]. Now, it can be verified by direct computation that the following vectors all belong to $\mathcal{E}^{\circ}$:
@@
With some more tedious arithmetic, we can also compute that
@@
see, e.g., [35]. Now, it can be verified by direct computation that the following vectors all belong to $P_{(2/3,1/3)}^{\circ}$:
@@
[…]
@@
In both cases, the certificate $y$ is easy to verify as long as the first two derivatives of the barrier function are easily computable (which in turn implies that membership in ${\mathcal{K}^{*}}$ is easily verifiable): one needs to first verify $y\in{\mathcal{K}^{*}}$ and compute the matrix $B$ or $H$ at $y$; the rest of the verification process amounts to simple linear algebra and verification of the membership of another vector in ${\mathcal{K}^{*}}$. In particular, many known LHSCBs have efficiently computable gradients and Hessians that are also rational at every rational point. In this case, the computed rational dual membership certificates can even be rigorously verified in rational arithmetic.
@@
Additionally, let $y_{b}$ be the unique vector in ${({\mathcal{K}^{*}})^{\circ}}$ that satisfies $-g(y_{b})=b$. (Recall Proposition 1.2.) Then $\mathcal{B}(b)$ and $\mathcal{H}(b)$ are full-dimensional subcones of ${({\mathcal{K}^{*}})^{\circ}}$ containing $y_{b}$.
@@
If a vector $y\in{\mathcal{K}^{*}}$ is a B-certificate for $b$, then $y$ is also an H-certificate for $b$. The converse, in general, does not hold.
@@ 3 A reconstruction formula for linear images of cones
We now turn our attention to our primary motivating examples, which are surjective linear images of high-dimensional cones. Suppose that
(9) $\mathcal{K}_{A}=\{Ax\,|\,x\in\mathcal{K}\},\text{ and therefore }\;{\mathcal{K}_{A}^{*}}=\{y\,|\,A^{\mathrm{T}}y\in{\mathcal{K}^{*}}\},$
the rows of $A$ are linearly independent, and assume that ${\mathcal{K}^{*}}$ has a non-empty interior that is the domain of a known, easily computable LHSCB. Even if $\mathcal{K}$ has a known LHSCB, its linear image $\mathcal{K}_{A}$ does not “inherit” an obvious barrier function; this is the main reason why it is difficult to optimize over $\mathcal{K}_{A}$, and why deciding feasibility in $\mathcal{K}_{A}$ is non-trivial even if deciding feasibility in $\mathcal{K}$ is easy. However, because the $A^{\mathrm{T}}$ map is injective, if ${\mathcal{K}_{A}^{*}}$ is full-dimensional, it inherits an easily computable LHSCB from ${\mathcal{K}^{*}}$: if $f$ is an LHSCB for ${\mathcal{K}^{*}}$, then
@@
is automatically an LHSCB with domain $({\mathcal{K}_{A}^{*}})^{\circ}$. (Recall Proposition 1.2.) The gradient and Hessian of $f_{A}$ can be expressed in closed form.
@@
(13) $x=B(A^{\mathrm{T}}y)A^{\mathrm{T}}\left(AB(A^{\mathrm{T}}y)A^{\mathrm{T}}\right)^{-1}b$
@@
The same statement holds for H-certificates if in (13) we replace every instance of the matrix $B$ with $H$.
@@
We detail the proof for B-certificates only; it can be translated to H-certificates word-for-word by, as suggested in the statement, substituting the Hessian $H$ in place of $B$ throughout the proof.
We first note that $x$ is well-defined: $B$ is well-defined and invertible, since $y\in(\mathcal{K}_{A}^{*})^{\circ}$ and $A^{\mathrm{T}}y\in({\mathcal{K}^{*}})^{\circ}$; additionally, $A$ is of full row-rank, which ensures that the matrix $\left(AB(A^{\mathrm{T}}y)A^{\mathrm{T}}\right)$ inverted in the formula is indeed invertible.
The first of our two claims, $Ax=b$, is immediate from basic arithmetic.
To show that $x\in\mathcal{K}$, first note that we can apply Eqs. (11)–(12) and Theorem 2.1 to the cone ${\mathcal{K}_{A}^{*}}$ and its LHSCB $f_{A}$ to see that
@@
and using Theorem 2.1 again, we obtain the desired relation
@@
This theorem gives a closed-form formula with which we can compute a suitable pre-image of $b\in\mathcal{K}_{A}$ that automatically belongs to $\mathcal{K}$ to make the verification of $b\in\mathcal{K}_{A}$ even simpler (as long as membership in $\mathcal{K}$ can be easily verified). In other words, the formula (13) converts the dual certificate $y$ into a conventional, “primal” certificate of $b$’s membership in $\mathcal{K}_{A}$. We note, however, that the relation $b\in\mathcal{K}_{A}$ is already rigorously certified by the relation $B_{A}(y)^{-1}b\in{\mathcal{K}_{A}^{*}}$, and the computation of the primal certificate $x$ is not necessary to produce or verify a dual certificate. On the contrary, dual certificates allow us to optimize directly over $\mathcal{K}_{A}$ and verify the feasibility of the computed (near-)optimal solutions without resorting to computing and verifying the membership of any vector in $\mathcal{K}$, which is useful whenever $\dim(\mathcal{K}_{A})\ll\dim(\mathcal{K})$. We detail the algorithmic aspects of these questions in Section 5.
@@
(14) $\min_{x}\left\{\left\|-g(A^{\mathrm{T}}y)-x\right\|_{B(A^{\mathrm{T}}y)^{-1}}\,\middle|\,Ax=b\right\},$
where $y\in{({\mathcal{K}^{*}})^{\circ}}$ is a given vector. Straightforward arithmetic involving (11)–(12) (without any reliance on Proposition 1.2) shows that if $\lambda$ denotes the Lagrange multipliers associated with the equality constraints, then the optimality conditions of (14) yield the optimal multipliers
$\lambda^{*}=B_{A}(y)^{-1}\left(-g_{A}(y)-b\right)$
@@
(15) $x^{*}=B(A^{\mathrm{T}}y)A^{\mathrm{T}}B_{A}(y)^{-1}b-\left(I-B(A^{\mathrm{T}}y)A^{\mathrm{T}}B_{A}(y)^{-1}A\right)g(A^{\mathrm{T}}y).$
Using (3), the second term simplifies to 00, after which the equation (15) reduces to (13).
The interpretation of (14) is fairly clear. The range of $-g$ is the primal cone $\mathcal{K}^{\circ}$, in which we wish $x$ to belong (recall the first claim of Proposition 1.2); thus, this problem seeks to find an $x$ satisfying $Ax=b$ that is close (in a suitable, $y$-dependent norm) to a given vector that is already in $\mathcal{K}^{\circ}$.
@@
$\left\|-g(A^{\mathrm{T}}y)-x^{*}\right\|_{B(A^{\mathrm{T}}y)^{-1}}=\left\|-g_{A}(y)-b\right\|_{B_{A}(y)^{-1}},$
and analogously, for H-certificates we may replace every occurrence of $B$ with $H$ in the previous calculations to obtain the optimal objective value $\left\|-g_{A}(y)-b\right\|_{y}^{*}.$ We can use these observations to paint a more geometric picture, and to derive easy-to-use sufficient conditions for dual certification. For this, we briefly recall the notion of Dikin ellipsoids.
@@
It is well-known that Dikin ellipsoids of radius $1$ stay entirely within their respective cones [34, Chapter 2]. Applying this to $\mathcal{K}_{A}$ and $\mathcal{K}$, we see that if

### v3

@@ Abstract
We discuss optimization problems over convex cones in which membership is difficult to verify directly. In the standard theory of duality, vectors in the dual cone $\mathcal{K}^{*}$} are associated with separating hyperplanes and interpreted as certificates of non-membership in the primal cone $\mathcal{K}$. Complementing this perspective, we develop easily verifiable certificates of membership in $\mathcal{K}$ given by vectors in $\mathcal{K}^{*}$}. Assuming that $\mathcal{K}^{*}$} admits an efficiently computable logarithmically homogeneous self-concordant barrier, every vector in the interior of $\mathcal{K}$ is associated with a full-dimensional cone of efficiently verifiable membership certificates. Consequently, rigorous certificates can be computed using numerical methods, including interior-point algorithms. The proposed framework is particularly well-suited to optimization over low-dimensional linear images of higher dimensional cones: we argue that these problems can be solved by optimizing directly over the (low-dimensional) dual cone, circumventing the customary lifting that introduces a large number of auxiliary variables. As an application, we derive a novel closed-form formula for computing exact primal feasible solutions from suitable dual feasible solutions; as the dual solutions approach optimality, the computed primal solutions do so as well. To illustrate the generality of our approach, we show that the new certification scheme is applicable to virtually every tractable subcone of nonnegative polynomials commonly used in polynomial optimization (such as SOS, SONC, SAGE and SDSOS polynomials, among others), facilitating the computation of rigorous nonnegativity certificates using numerical algorithms.
@@ 1 Introduction
for every $z\in\mathcal{K}\setminus\{0\}$ [14, Chapter 6]. This vector $y$ can be interpreted as a non-membership certificate, that is, a proof that $b$ does not belong to $\mathcal{K}$. By definition, $y$ is an element of the interior of the dual cone of $\mathcal{K}$,
@@
this gives rise to the standard duality theory of convex (conic) programming. In this paper we show that with the help of self-concordance theory, membership (rather than non-membership) in the primal cone can be similarly certified by simply exhibiting suitable vectors from the dual cone. These certificates are not unique; on the contrary, each element of the interior of $\mathcal{K}$ has a full-dimensional cone of membership certificates. This facilitates their computation using numerical methods and has important theoretical consequences.
Our first main contribution is that whenever $\mathcal{K}^{*}$} is a hyperbolicity cone (see [12, 19] for the definition and standard examples), membership in $\mathcal{K}$ can be certified in a manner similar to the dual sums-of-squares certificates of [8] (we call these H-certificates in Definition 2.4), but the generalization does not go any further (Example 2.3). Then we present a new certification scheme (termed B-certificates) that is equally efficient for the same cones as H-certificates, but can also be applied to all cones $\mathcal{K}$ as long as $({\mathcal{K}^{*}})^{\circ}$} is the domain of an efficiently computable logarithmically homogeneous self-concordant barrier function (Theorem 2.6).
@@
In Section 3, we demonstrate that our dual membership certificates overcome both of these issues. First, because the cone of certificates for every vector is full-dimensional, numerically computed certificates are automatically exact, rigorous certificates; and since they are just elements of $\mathcal{K}_{A}^{*}$, they are not higher dimensional than the certified vector. Second, the certification schemes come with a reconstruction formula for cones of the form (1) that allows the efficient computation of a primal membership certificate $x\in\mathcal{K}$ that by definition satisfies $b=Ax$ exactly. (Theorem 3.2.) For most cones $\mathcal{K}$ that commonly arise in conic optimization problems, the formula automatically yields rational primal certificates $x$ (as long as $A$ is a rational matrix), which can also be verified efficiently in rational arithmetic.
In Section 6, we show how dual certificates can be used to compute, with a similar closed-form formula, exact primal feasible solutions of conic optimization problems in standard form from suitable dual feasible solutions, with a guarantee that the closer the dual solutions are to optimality, the closer the computed primal feasible solutions will be to optimality. With this formula, most short-step, path-following interior-point algorithms, including dual-only methods, can be “retrofitted” to compute primal-dual feasible pairs with a guaranteed low duality gap with minimal additional computational cost in each iteration.
In Section 7, we outline the application of these ideas in polynomial optimization. The most commonly used inner approximations of the cones of nonnegative polynomials (including sums-of-squares (SOS) polynomials, sums of nonnegative circuit (SONC) polynomials, sums of AM-GM exponentials (aka. SAGE functions), SDSOS polynomials and others) are all cones of the form (1) with an easily computable matrix $A\in\mathbb{Q}^{m\times n}$, usually with $m\ll n$, and a (possibly non-symmetric) cone $\mathcal{K}$ for which $\mathcal{K}^{*}$} admits a well-known logarithmically homogeneous self-concordant barrier function. Therefore, the problem of rigorously certifying the nonnegativity of a polynomial using either of these cones amounts to the exact solution of a feasibility problem in $\mathcal{K}_{A}$, as discussed above. When $\mathcal{K}_{A}$ is any of the aforementioned cones of polynomials, the dual certificates introduced in this paper can be interpreted as easily verifiable rigorous polynomial nonnegativity certificates that can be computed numerically without constructing an explicit “primal” representation (such as an explicit representation of an SOS polynomial as a sum of squared polynomials). Should such a representation be necessary, the reconstruction formula can be used to compute an exact representation in closed form using a numerically computed dual certificate.
@@
The results in this paper are broad generalizations of prior work of Davis and Papp [8] on weighted sums-of-squares (WSOS) cones, which are convex cones instrumental in global polynomial optimization. Membership in these cones can be cast as a (feasibility) semidefinite program (SDP) [32]; every feasible solution to this SDP is an obvious membership certificate (a Gram matrix of the SOS polynomial). However, the number of variables in this SDP is at least an order of magnitude larger than the dimension of the cone, and in the presence of a large number of weights, the ratio of the dimensions of these two cones can be arbitrarily large. Instead, the certificates introduced in [8] (simply termed there as dual certificates) are elements of the dual WSOS cone, and are thus low-dimensional, and can be more efficiently computed using non-symmetric cone programming methods than solving the conventional SDP. The theory presented in the present paper is a generalization of these earlier results, and the algorithms discussed here are simultaneously more general and more efficient (owing to their lower iteration complexity) than the algorithms presented in the earlier work.
@@
    $f(x)\to\infty$ as $x\to\operatorname{bd}(\mathcal{C})$.
@@
In the rest of the paper, we will continue to use $g$ and $H$ to denote the gradient and the Hessian of the SCB $f$.
@@
Proposition 1.2 below summarizes a few properties of self-concordant barrier functions used throughout the paper. The proofs of these statements can be found, for example, in [36, Chapters 2–3] and [26, Chapter 5].
@@
Suppose $f:\mathcal{C}^{\circ}\to\mathbb{R}$ is an SCB for some convex cone $\mathcal{C}\subseteq\mathbb{R}^{n}$. Then the following hold:
@@
    The gradient map $g:\mathcal{C}^{\circ}\to\mathbb{R}^{n}$ is a bijection between $\mathcal{C}^{\circ}$ and $(\mathcal{C}^{*})^{\circ}$. That is, for every $b\in(\mathcal{C}^{*})^{\circ}$ there exists a unique $y_{b}\in\mathcal{C}^{\circ}$ that satisfies $g(y_{b})=b$.
@@
    If $A:\mathbb{R}^{m}\to\mathbb{R}^{n}$ is an injective linear operator and $\mathcal{A}\overset{\mathrm{def}}{=}\left\{x\in\mathbb{R}^{m}\,|\,A(x)\in\mathcal{C}^{\circ}\right\}$ is non-empty, then $x\mapsto f(A(x))$ is a $\nu$-LHSCB with domain $\mathcal{A}$.
@@ 2 Membership certificates from the dual cone
Throughout this paper, we shall use the following assumptions and notational shortcuts: we assume that $\mathcal{K}$ is a proper convex cone, i.e., it is convex, closed, full-dimensional (has a non-empty interior), and pointed; in addition, we assume that the dual cone $\mathcal{K}^{*}$} admits a $\nu$-LHSCB $f$, and we denote the gradient and Hessian of $f$ at $y\in{({\mathcal{K}^{*}})^{\circ}}$ by $g(y)$ and $H(y)$, respectively. Finally, we define
@@
The membership certification schemes introduced in this section are based on the following facts. Note that the first half of the next theorem holds under the less restrictive assumption that $f$ is an SCB, rather than an LHSCB.
@@
Let $\mathcal{K}$ be a proper convex cone whose dual cone $\mathcal{K}^{*}$} admits an SCB $f$. Then, with the above notation, for every $y\in{({\mathcal{K}^{*}})^{\circ}}$ and $z\in{\mathcal{K}^{*}}$,
@@
If $\mathcal{K}^{*}$} is the hyperbolicity cone of the hyperbolic polynomial $p$ and $f$ is the logarithmic barrier $\ln(p(\cdot))$, then we also have
@@
The second claim is Güler’s result from [13, Thm. 6.2].
The assumption of hyperbolicity in the second half of Theorem 2.1 is necessary; the relation (5) does not hold for every cone, as the following examples show.
@@
$\mathcal{E}\overset{\mathrm{def}}{=}\operatorname{cl}\left(\left\{(x_{1},x_{2},x_{3})\in\mathbb{R}_{+}^{2}\times\mathbb{R}\,|\,x_{2}>0,x_{1}\geq x_{2}e^{x_{3}/x_{2}}\right\}\right),$
@@
see, e.g., [25, Eq. (4.5)]. Now, it can be verified by direct computation that the following vectors all belong to $\mathcal{E}^{\circ}$:
@@
With some rather tedious arithmetic we can also compute that
@@
see, e.g., [37]. Now, it can be verified by direct computation that the following vectors all belong to $\mathcal{P}_{(2/3,1/3)}^{\circ}$:
@@
$w^{\mathrm{T}}H(y)z=\frac{12\,871-443\,620\sqrt[3]{10}+195\,500\sqrt[3]{100}}{60\left(1-10\sqrt[3]{10}\right)^{2}}<-1.399,$
@@
In both cases, the certificate $y$ is easy to verify as long as the first two derivatives of the barrier function are easily computable (which in turn implies that membership in $\mathcal{K}^{*}$} is easily verifiable): one needs to first verify $y\in{({\mathcal{K}^{*}})^{\circ}}$ and compute the matrix $B$ or $H$ at $y$; the rest of the verification process amounts to simple linear algebra and verification of the membership of another vector in $\mathcal{K}^{*}$}. In particular, many known LHSCBs have efficiently computable gradients and Hessians that are also rational at every rational point. In this case, the computed rational dual membership certificates can even be rigorously verified in rational arithmetic.
@@
Additionally, let $y_{b}$ be the unique vector in $({\mathcal{K}^{*}})^{\circ}$} that satisfies $g(y_{b})=b$. (Recall Proposition 1.2.) Then $\mathcal{B}(b)$ and $\mathcal{H}(b)$ are full-dimensional subcones of $({\mathcal{K}^{*}})^{\circ}$} containing $y_{b}$.
@@
If a vector $y\in{({\mathcal{K}^{*}})^{\circ}}$ is a B-certificate for $b$, then $y$ is also an H-certificate for $b$. The converse, in general, does not hold.
@@ 3 A reconstruction formula for linear images of cones
We now turn our attention to our primary motivating examples, which are surjective linear images of high-dimensional cones. Let
(9) $\mathcal{K}_{A}=\{Ax\,|\,x\in\mathcal{K}\}\text{ and therefore }\;{\mathcal{K}_{A}^{*}}=\{y\,|\,A^{\mathrm{T}}y\in{\mathcal{K}^{*}}\},$
where $A\in\mathbb{R}^{m\times n}$, the rows of $A$ are linearly independent, and assume that $\mathcal{K}^{*}$} has a non-empty interior that is the domain of a known, easily computable LHSCB. Even if $\mathcal{K}$ has a known LHSCB, its linear image $\mathcal{K}_{A}$ does not “inherit” an obvious barrier function; this is the main reason why it is difficult to optimize over $\mathcal{K}_{A}$ and why deciding feasibility in $\mathcal{K}_{A}$ is non-trivial even if deciding feasibility in $\mathcal{K}$ is easy. However, under mild conditions, $\mathcal{K}_{A}^{*}$} inherits an easily computable LHSCB from $\mathcal{K}^{*}$}. More precisely, using Proposition 1.2, if $f$ is an LHSCB for $\mathcal{K}^{*}$}, $A^{\mathrm{T}}$ is injective, and $\mathcal{K}_{A}^{*}$} is full-dimensional, then the function $f_{A}:({\mathcal{K}_{A}^{*}})^{\circ}\to\mathbb{R}$ defined by
@@
is automatically an LHSCB for $\mathcal{K}_{A}^{*}$}. The gradient and Hessian of $f_{A}$ can be expressed in closed form.
@@
Our next theorem gives a closed-form formula with which we can compute a suitable pre-image of $b\in\mathcal{K}_{A}$ that automatically belongs to $\mathcal{K}$. This makes the verification of $b\in\mathcal{K}_{A}$ even simpler (as long as membership in $\mathcal{K}$ can be easily verified).
@@
(13) $x=H(A^{\mathrm{T}}y)A^{\mathrm{T}}\!\left(AH(A^{\mathrm{T}}y)A^{\mathrm{T}}\right)^{-1}b$
@@
We first note that $x$ is well-defined: $H(A^{\mathrm{T}}y)$ is well-defined and invertible, since $y\in(\mathcal{K}_{A}^{*})^{\circ}$ and $A^{\mathrm{T}}y\in({\mathcal{K}^{*}})^{\circ}$; additionally, $A$ is of full row-rank, which ensures that the matrix $\left(AH(A^{\mathrm{T}}y)A^{\mathrm{T}}\right)$ inverted in the formula is indeed invertible.
The first of our two claims, $Ax=b$, is immediate from simple matrix arithmetic.
To show that $x\in\mathcal{K}$, first note $B_{A}(y)$ is also invertible. Then we can apply Eqs. (11)–(12) and Theorem 2.1 to the cone $\mathcal{K}_{A}^{*}$} and its LHSCB $f_{A}$ and see that
@@
and using Theorem 2.1 again, we obtain the relation
@@
Finally, note that
$B(A^{\mathrm{T}}y)A^{\mathrm{T}}(AB(A^{\mathrm{T}}y)A^{\mathrm{T}})^{-1}=H(A^{\mathrm{T}}y)A^{\mathrm{T}}(AH(A^{\mathrm{T}}y)A^{\mathrm{T}})^{-1}$
by repeated application of (3) in Proposition 1.2.
Note that Theorem 3.2 does not require a separate formula for H-certificates, since (for cones that admit H-certificates) every H-certificate is automatically a B-certificate.
The formula (13) converts the dual certificate $y$ into a conventional, “primal” certificate of $b$’s membership in $\mathcal{K}_{A}$. We note, however, that the relation $b\in\mathcal{K}_{A}$ is already rigorously certified by the relation $B_{A}(y)^{-1}b\in{\mathcal{K}_{A}^{*}}$, and the computation of the primal certificate $x$ is not necessary to produce or verify a dual certificate. On the contrary, dual certificates allow us to optimize directly over $\mathcal{K}_{A}$ and verify the feasibility of the computed (near-)optimal solutions without resorting to computing and verifying the membership of any vector in $\mathcal{K}$, which is useful whenever $\dim(\mathcal{K}_{A})\ll\dim(\mathcal{K})$. We detail the algorithmic aspects of these questions in Section 5.
@@
(14) $\min_{x}\left\{\left\|-g(A^{\mathrm{T}}y)-x\right\|_{H(A^{\mathrm{T}}y)^{-1}}\,\middle|\,Ax=b\right\},$
where $y\in{({\mathcal{K}^{*}})^{\circ}}$ is a given vector. Straightforward arithmetic involving (11) (without any reliance on Proposition 1.2) shows that if $\lambda$ denotes the Lagrange multipliers associated with the equality constraints, then the optimality conditions of (14) yield the optimal multipliers
$\lambda^{*}=y-H_{A}(y)^{-1}b$
@@
(15) $x^{*}=H(A^{\mathrm{T}}y)A^{\mathrm{T}}H_{A}(y)^{-1}b,$
exactly as in the reconstruction formula (13).
The interpretation of (14) is fairly clear. The range of g-g is $\mathcal{K}^{\circ}$, in which we wish $x$ to belong (recall the first claim of Proposition 1.2); thus, this problem seeks to find an $x$ satisfying $Ax=b$ that is close (in a suitable, $y$-dependent norm) to a given vector that is already in $\mathcal{K}^{\circ}$.
@@
$\left\|-g(A^{\mathrm{T}}y)-x^{*}\right\|_{H(A^{\mathrm{T}}y)^{-1}}=\left\|-g_{A}(y)-b\right\|_{y}^{*},$
keeping in mind that the dual local norm $\|\cdot\|^{*}_{y}$ now corresponds to the barrier $f_{A}$.
We can use these observations to paint a more geometric picture, and to derive easy-to-use sufficient conditions for dual certification. For this, we briefly recall the notion of Dikin ellipsoids.
@@
It is well-known that Dikin ellipsoids of radius $1$ stay entirely within their respective cones [36, Chapter 2]. Applying this to $\mathcal{K}_{A}$ and $\mathcal{K}$, we see that if